OpenStudy (anonymous):

How many solutions does the system of equations have? y - 7x = -14 and 7y --49x = -2

OpenStudy (anonymous):

y-7x=-14 7y+49x=-2 add 7x subtract 49x y=7x-14 7y=-49x-2 divide by 7 y=7x-14 y=-7x-2 different slopes so one

OpenStudy (anonymous):

Thank you again! No more questions! THanks!

OpenStudy (anonymous):

no problem